40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1040     Entry time: Mon Aug 27 14:24:06 2012
Author: tara 
Type: Notes 
Category: Schedule 
Subject: plan for upgraded thermal noise probe experiment 

 

 The 1.45" cavities have arrived. So I think it is a good time I layout the plan for the next phase of the experiment:

IMG_1579.JPG

 

To Do:

  • Design optic layout for 2-laser setup
  • Try assemble the cavities with our SiO2/Ta2O5 mirrors
  • modify seismic stack / design for holding the cavities
  • getting AlAs/GaAs coating on the substrates
  • find substrate with AdvLIGO coating
  • Choose resonant frequency for 2nd PMC, and 2nd cavity

 

2_laser_setup.png

fig2: proposed new setup with 2 laser sources

Since we want to upgrade the current setup to 2-laser setup, we need to find:

  • another NPRO
  • another PMC
  • broad band EOM, and two resonant EOMs (one for locking PMC, one for locking to the cavity)
  • RFPD, servo, oscillator for locking 2nd PMC to NPRO
  • RFPD for 2nd cavity (with the correct resonant frequency)
  • second oscillator for locking NPRO to the cavity (we can reuse the marconi for AOM)
  • another servo for PMC (I think there is one spare that we used for driving a PZT, I'll check)
ELOG V3.1.3-